Lemon Inc. has the following information regarding its preferred and common stock:

Preferred stock, $30 par, 15% cumulative; 300,000 shares authorized; 180,000 shares issued and outstanding
Common stock, $2 par; 2,500,000 shares authorized; 1,800,000 shares issued; 1,600,000 outstanding

As of December 31, 2019, Lemon was 3 years in arrears on its dividends. During 2020, Lemon declared and paid dividends. As a result, the common stockholders received dividends of $0.80 per share.

CHART OF ACCOUNTS
Lemon Inc.
General Ledger
ASSETS
111 Cash
112 Petty Cash
121 Accounts Receivable
123 Allowance for Doubtful Accounts
124 Notes Receivable
125 Supplies
126 Prepaid Insurance
127 Prepaid Rent
128 Inventory
129 Interest Receivable
130 Investment
131 Land
132 Buildings
133 Equipment
134 Furniture
135 Trucks
139 Accumulated Depreciation
141 Leased Assets
LIABILITIES
211 Accounts Payable
212 Utilities Payable
213 Insurance Payable
214 Wages Payable
215 Rent Payable
221 Interest Payable
223 Income Taxes Payable
231 Dividends Payable
232 Unearned Sales Revenue
233 Unearned Rent Revenue
241 Bonds Payable
251 Notes Payable
EQUITY
311 Common Stock
312 Additional Paid-In Capital-Common Stock
313 Preferred Stock
314 Additional Paid-In Capital-Preferred Stock
315 Treasury Stock
316 Additional Paid-In Capital-Treasury Stock
321 Retained Earnings
331 Dividends
REVENUE
411 Sales Revenue
412 Service Revenue
413 Interest Income
414 Rent Revenue

Required:

1. What was the total amount of dividends declared and paid?

2. What journal entry was made at the date of declaration?
Answer #1
1
Arrear Preferred Dividends 2430000 =180000*30*15%*3
Current Preferred Dividends 810000 =180000*30*15%
Common stock Dividends 1280000 =1600000*0.80
Total amount of dividends declared and paid 4520000
2
Debit Credit Assets Liabilities Equity
Dividends 4520000
      Dividends Payable 4520000